OpenStudy (anonymous):

y varies jointly with p and q. When p = 4 and q = 5, y = 60. Find y when p = 5 and q = 8. PLEASE Help :(

OpenStudy (anonymous):

y = k p q 60 = k (5)(4) 60= 20 k k=3

OpenStudy (anonymous):

y=k(pq)

OpenStudy (anonymous):

y= 3 p q y = 3 (5)(8)

OpenStudy (anonymous):

Do the way imran is doing..........

OpenStudy (anonymous):

i dont know how to do itttt, i just dont get it

OpenStudy (anonymous):

y=k(pq) Now when p=4,q=5,y=60so 60=k(4*5) k=3 so the equation becomes y=3pq now put p=5 and q=8 y=3*5*8=120